OpenStudy (anonymous):

What is the point-slope equation of the line with slope -1/3 that goes through the point (4, -5)?

OpenStudy (karatechopper):

y-y1=m(x-x1) That is the Point Slope Formula^

OpenStudy (karatechopper):

Do you know what the slope is?

OpenStudy (anonymous):

nope

OpenStudy (karatechopper):

Nvm, its in there. Okay. So what you are going to do basically is put all the numbers in and simplify.

OpenStudy (karatechopper):

y1- Y coordinate x1- X coordinate m- Slope

OpenStudy (karatechopper):

In this case y1=-5 x1=4 m= -1/3

OpenStudy (karatechopper):

Do you Agree?

OpenStudy (karatechopper):

@ChinaGurl you there..?

OpenStudy (anonymous):

yes! so its A!

OpenStudy (karatechopper):

Uh. Whats A?

OpenStudy (anonymous):

y - 5 + -1/3 (x + 4)

OpenStudy (anonymous):

@karatechopper

OpenStudy (karatechopper):

Actually...no. You kind of messed up there. Lets backtrack. Soo..you put the numbers in.. y - (- 5) = -1/3 (x + 4) Then simplify. So it would be.. y+5= -1/3 (x +4)

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

the weird part is your answer is not part of the answer choices! @karatechopper

OpenStudy (karatechopper):

Can you show me what the question looked like?

OpenStudy (anonymous):

yes